# Optimal navigation in two-dimensional regular and turbulent flows

### v1

@@ VI.2 Example 1: Snapshot of 2D Turbulence
We begin by examining the navigation problem in a single snapshot of the 2d turbulent flow. The training hyperparameters for both the Q-learning and one-step actor-critic algorithms are identical to those used in Section LABEL:sec:RL_2d_turb_snap. For all considered filtering thresholds, $k_{max}=\{10,5,3,2\}$, the learning curves for both methods reach a clear plateau, indicating convergence toward effective control policies.
Representative trajectories obtained from these learned policies – evaluated in the full (unfiltered) velocity field – are shown in Fig. LABEL:fig:15. For comparison, we also display the optimal control trajectory computed in Section LABEL:sec:OPC_D. As illustrated, the learned policies remain robust even when the agent is provided only with coarse-grained flow information: for $k_{max}\geq 3$, both algorithms reliably guide the agent to the target along paths that remain reasonably close to the optimal solution. Consistent with our earlier findings, the one-step actor-critic agent performs slightly better than the Q-learning agent. However, the difference in their mean travel times does not exceed $10\%$, and overall both RL methods succeed in solving the task with zero failure rate, see Table LABEL:tab:1.
When the filtering becomes more severe, at $k_{max}=2$, the situation changes qualitatively. The Q-learning agent is no longer able to reach the target at all. The actor-critic agent performs somewhat better, achieving success in roughly $20\%$ of trials, but its average travel time nearly doubles compared with the optimal control trajectory. In summary, removing velocity modes with wavenumbers $k>2$ significantly distorts the flow structure, to the point where the learned policies become practically useless.
Table $\labeltab:1$Example of the 2d turbulence snapshot: Mean travel time $\langle T\rangle$ of successful target-reaching attempts and failure rates across different filtering thresholds $k_{max}$.
Q-learning One-Step Actor-Critic
$k_{max}$ $\langle T\rangle$ failure rate $\langle T\rangle$ failure rate
$\infty$ $0.96$ 0 $0.87$ 0
$10$ $0.93$ 0 $0.87$ 0
55 $0.94$ 0 $0.91$ 0
33 $0.91$ 0 $0.88$ 0
22 11 $1.65$ $0.77$
@@ VI.3 Example 2: 2D Turbulence
The next example focuses on a navigation problem in time-dependent 2d turbulent flow. The training hyperparameters for both the Q-learning and one-step actor-critic methods are selected to be identical to those used earlier in Section LABEL:sec:RL_2d_turb_time. For all filtering thresholds $k_{max}=\{10,5,3,2\}$, we achieved convergence of the RL algorithms toward effective control policies, as confirmed by the learning curves reaching a plateau.
Figure LABEL:fig:16 shows the agent trajectories corresponding to the learned policies for different filtering scales $k_{max}$ of velocity field. For reference, we also display the trajectory obtained from the optimal control theory in Section LABEL:sec:OPC_E. Consistent with our earlier findings, the RL algorithms converge to a trajectory distinct from the optimal control path. This RL-derived solution remains robust with respect to the removal of small-scale velocity fluctuations with $k\gtrsim k_{f}$. When the filtering scale $k_{max}$ is reduced further, the agent trajectories become increasingly irregular. Moreover, in rare cases, the agent fails to reach the target, and for those attempts that do succeed, the average travel time increases, see Table LABEL:tab:2. When comparing the Q-learning and one-step actor-critic methods, the action-value approach seems to performs slightly better at low $k_{max}$.
At a qualitative level, the example of time-dependent 2d turbulent flow considered here behaves similarly to the previously studied example of 2d turbulent snapshot. The trajectories produced by the RL algorithms exhibit robustness to the absence of small-scale velocity field information; however, once the filtering scale approaches the characteristic size of the flow structures that guide the agent’s path selection, the performance of the methods deteriorates noticeably. For 2d turbulence, the critical filtering scale $k_{max}$ lies within the inverse energy cascade range.
Table $\labeltab:2$Example of the time-dependent 2d turbulent flow: Mean travel time $\langle T\rangle$ of successful target-reaching attempts and failure rates across different filtering thresholds $k_{max}$.
Q-learning One-Step Actor-Critic
$k_{max}$ $\langle T\rangle$ failure rate $\langle T\rangle$ failure rate
$\infty$ $0.94$ 0 $0.97$ 0
$10$ $0.93$ 0 $0.94$ 0
55 $0.99$ 0 $0.97$ 0
33 $0.99$ $0.05$ $1.27$ $0.01$
22 $1.34$ $0.04$ $3.01$ $0.27$
[…]
@@ VII Conclusion
In this work, we conducted a systematic investigation of optimal navigation in two-dimensional flows of increasing dynamical complexity, ranging from steady vortical fields to fully developed turbulence in the inverse energy cascade regime. Using classical optimal control theory, we demonstrated that time-minimizing trajectories can be computed accurately even in complex environments; however, their practical applicability becomes limited in complex flows due to strong sensitivity to initial conditions and the instability of the resulting optimal paths. These issues highlight the fundamental challenges associated with implementing pre-computed optimal-control strategies in realistic, chaotic settings where small perturbations and incomplete knowledge of the environment are unavoidable.
@@
To address these limitations, we employed RL as a model-free alternative capable of adapting to complex flow structures without relying on explicit knowledge of the governing dynamics. By comparing two canonical classes of RL algorithms – action-value (Q-learning) and policy gradient (one-step actor-critic) methods – we showed that both approaches reliably produce robust navigation strategies across all considered flow configurations. While the RL-derived trajectories do not always coincide with the optimal-control solutions, they remain stable and consistently achieve successful target acquisition, even in fully time-dependent turbulence.
@@
A key result of our study is the strong generalization capability of RL agents trained on coarse-grained velocity fields. Agents exposed only to large-scale flow information were nevertheless able to navigate efficiently within the full (unfiltered) turbulent environment. This robustness to missing small-scale data is particularly promising for real-world applications, where environmental forecasts are typically low-resolution and subject to uncertainty.
@@
The framework introduced here also offers a flexible platform for future studies. The use of deep RL, where action-value functions and policies are represented by neural networks, would eliminate the need for state-action discretization and enable navigation in more complex geometries or higher-dimensional velocity fields. Next, the objective functional itself can be generalized: besides minimizing travel time, one may incorporate additional criteria such as energy expenditure, fuel consumption, or multi-objective trade-offs between efficiency and safety [buzzicotti2020optimal]. Another important direction of research can be related to more realistic agent models. Incorporating inertia, turning-rate limitations and stochastic forcing representing environmental noise would further bridge the gap between theory and practical deployment [colabrese2017flow, alageshan2020machine, piro2022optimal].
@@
From a methodological viewpoint, the coexistence of ground truth optimal-control solutions and RL-based strategies within the same testbed creates a rare opportunity. It allows quantitative comparison between exact minimum-time trajectories and learned policies in data-driven control. Because the navigation problem is physically transparent, the framework is also well suited for use in teaching, providing an accessible setting in which students can explore nonlinear dynamics, turbulent transport, and modern machine-learning techniques within a unified context.
@@
Overall, our findings demonstrate that RL provides a powerful and resilient framework for navigation in complex flow environments, complementing classical optimal control theory and overcoming its primary limitations in chaotic regimes. These results open the door to developing practical autonomous navigation strategies for micro-robots, underwater vehicles, and other active agents operating in realistic oceanic and atmospheric flows.
@@ Acknowledgements.
The work was supported by the Ministry of Science and Higher Education of the Russian Federation and by the Basic research program of HSE University.
@@ Appendix A Incremental update
[…]
sec:appA
Assume for a moment that we are interested in estimating the mean value of a random number generator. At each step, we obtain a single sample, and just before the $n$-th step, our best estimate of the mean is given by
[…]
@@
where the target represents a potentially noisy indicator of the true value toward which the estimate should move, and $\alpha$ is the step-size or learning rate that controls how strongly the estimate is updated. In the present case, the target corresponds to the new sample $r_{n}$, and the learning rate is $\alpha=1/n$. Because this learning rate decays over time, each update becomes smaller, which guarantees convergence of $Q_{n}$ to the true mean as $n\to\infty$ in accordance with the law of large numbers, see Eq. (LABEL:eq:A1).
@@ Appendix B Policy Gradient Theorem
$label{eq:b1}\nabla v_{\pi}(s)=\sum_{a}\left[\nabla\pi(a|s)q_{\pi}(s,a)+\pi(a|s)\nabla q_{\pi}(s,a)\right].$ (57)
@@
$label{eq:b3}\nabla q_{\pi}(s,a)=\sum_{s^{\prime}}p(s^{\prime}|s,a)\nabla v_{\pi}(s^{\prime}),$ (59)
where $p(s^{\prime}|s,a)=\sum_{r}p(s^{\prime},r|s,a)$ denotes the marginal transition probability. Substituting Eq. (LABEL:eq:b3) into Eq. (LABEL:eq:b1) leads to a recursive form,
[…]
The term involving $\nabla v_{\pi}(s^{\prime})$ can itself be expanded using Eq. (LABEL:eq:b4), and this substitution can be applied recursively to capture contributions from all future states along a trajectory generated by the policy. After the first substitution, we obtain

### v2

@@ VI.2 Example 1: Snapshot of 2D Turbulence
We first study the navigation problem in a single snapshot of the 2d turbulent flow. For all filtering scales $k_{max}$, the training hyperparameters for both the Q-learning and one-step actor-critic algorithms are identical to those employed in Section V.7. Representative trajectories of agents trained in coarse-grained flows are illustrated in Fig. 15. The analysis of the success probability $\langle p\rangle$ and the mean travel time $\langle T\rangle$ is provided in Fig. 16.
Figure 16: The success rate (top) and the mean travel time (bottom) vs. normalized filtering scale $k_{max}/k_{f}$ for the 2d turbulence snapshot example. Error bars indicate the standard deviation.
The learned policies remain reliable when the filtering scale satisfies $k_{max}\geq 3$. In this regime, both RL algorithms guide the agent toward the target along trajectories that closely resemble the optimal control solution, see Fig. 15. Consistent with our earlier observations, the one-step actor-critic algorithm outperforms Q-learning for the 2d turbulence snapshot example. It reaches the target with unit success probability, $\langle p\rangle=1$, across ten independent training runs and yields a smaller mean travel time $\langle T\rangle$, see Fig. 16. By contrast, the Q-learning agent is less robust. For $k_{max}=5$ and $k_{max}=3$, training fails in approximately one out of ten independent runs, in which case the agent is unable to find a path to the target. Even in successful runs, the mean travel time is larger than that of the actor-critic agent, and the variability across different trained models is higher. Notably, the stochasticity of the training process increases as $k_{max}$ decreases, reflecting the growing loss of information in the filtered velocity field.
When the filtering becomes more severe, at $k_{max}=2$, the behavior changes qualitatively. The Q-learning agent completely fails to reach the target, resulting in a vanishing success probability, $\langle p\rangle=0$. The actor-critic agent performs better, achieving success in roughly $50\%$ of the trials; however, its mean travel time nearly doubles compared to models trained on velocity fields that retain more spectral content. Consequently, the agent trajectories deviate markedly from the optimal control solution, see the right panel in Fig. 15. Thus, the removal of velocity modes with wavenumbers $k>2$ leads to a substantial distortion of the flow structure. At this level of coarse graining, the available information becomes insufficient for effective learning, rendering the resulting control policies largely impractical.
Figure 17: Trajectories produced by the Q-learning (red) and one-step actor-critic (blue) agents are shown together with the optimal control trajectory (white) for the time-dependent 2d turbulent flow. The background depicts the coarse-grained velocity field, retaining only modes with wavenumbers $k\leq k_{max}$, which was used during training. All trajectories, however, were evaluated in the full (unfiltered) velocity field.
@@ VI.3 Example 2: 2D Turbulence
The next example focuses on a navigation problem in time-dependent 2d turbulent flow. For all filtering scales $k_{max}$, the training hyperparameters for both the Q-learning and one-step actor-critic methods are selected to be identical to those used earlier in Section V.8. Representative trajectories of agents trained in coarse-grained flows are shown in Fig. 17. The corresponding statistical analysis of the success probability $\langle p\rangle$ and the mean travel time $\langle T\rangle$ is presented in Fig. 18.
Figure 18: The success rate (top) and the mean travel time (bottom) vs. normalized filtering scale $k_{max}/k_{f}$ for the time-dependent 2d turbulent flow. Error bars indicate the standard deviation.
The obtained results indicate that both algorithms are capable of solving the navigation task for a moderate level of coarse-graining $k_{max}\geq 3$. The one-step actor-critic approach exhibits a high success probability $\langle p\rangle\approx 1$. This is because the trajectories discovered by the agent exploit a strong background current directed toward the target, which does not require highly coordinated control actions. In contrast, the Q-learning agent follows trajectories that pass through a region of weak external flow, where successful navigation demands a more coordinated sequence of actions. As a result, the probability of reaching the target is lower, but the successful trajectories are closer to the theoretical solution predicted by optimal control theory and yield a smaller mean travel time, see Fig. 18.
Upon further increasing the coarse-graining scale to $k_{max}=2$, the success probability of the actor-critic algorithm drops sharply. This degradation is attributed to the strong fluctuations arising from unresolved velocity-field scales. The Q-learning algorithm, on the other hand, exhibits the opposite behavior: its probability of successfully reaching the target increases and approaches $\langle p\rangle\approx 1$. This behavior appears to result from a fortuitous alignment of circumstances: the policy learned in the coarse-grained flow is more robust to external noise and leads to a shorter mean travel time. We believe that this counterintuitive outcome is specific to the particular example considered here (the choice of initial and target locations). In more general settings, one would expect a reduced success probability and increased travel times, consistent with the trend observed when varying $k_{max}$ from $10$ down to $3$.
Summarizing the results of this and the preceding example, we observe that the trajectories produced by the RL algorithms exhibit a degree of robustness to the absence of small-scale velocity-field information. However, once the filtering scale approaches the characteristic size of the flow structures that govern the agent’s route selection, the performance of RL methods deteriorates markedly. Our results suggest that, for two-dimensional turbulence, the critical filtering scale $k_{max}$ is smaller than the characteristic forcing wavenumber $k_{f}$ and lies within the inverse energy cascade range.
@@ VII Conclusion
In this work, we conducted a systematic investigation of optimal navigation in two-dimensional flows of increasing dynamical complexity, ranging from steady solid-body rotation to fully developed turbulence in the inverse energy cascade regime. Using optimal control theory, we demonstrated that time-minimizing trajectories can be computed accurately even in complex environments; however, their practical applicability becomes limited in complex flows due to strong sensitivity to initial conditions and the instability of the resulting optimal paths. The presence of an additional degree of freedom associated with the steering direction makes the active system more susceptible to chaotic behavior than a passive scalar. In particular, this manifests as an exponential divergence of initially nearby trajectories of an active agent in some two-dimensional incompressible stationary velocity fields.
@@
To overcome the instability of optimal trajectories and identify stable behavioral strategies for agents, we employed RL as an alternative capable of adapting to complex flow structures. By comparing two canonical classes of RL algorithms – action-value (Q-learning) and policy gradient (one-step actor-critic) methods – we showed that both approaches produce robust navigation strategies across all considered flow configurations. The resulting agent trajectories do not always coincide for the two RL algorithms, and our analysis indicates that the Q-learning typically produces outcomes that more closely resemble the theoretical optimal control solutions.
@@
A key result of our study is the strong generalization capability of RL agents trained on coarse-grained turbulent velocity fields. Agents exposed only to large-scale flow information were nevertheless able to navigate efficiently within the full turbulent environment. This robustness to missing small-scale data is particularly promising for real-world applications, where environmental forecasts are typically low-resolution and subject to uncertainty. Importantly, the success of RL in this setting appears to stem less from algorithmic complexity than from the physical structure of the flow itself.
@@
The framework presented here also offers a flexible platform for future studies. The use of deep RL, where action-value functions and policies are represented by neural networks, would eliminate the need for state-action discretization and enable navigation in more complex settings or higher-dimensional velocity fields. Next, the objective functional itself can be generalized: besides minimizing travel time, one may incorporate additional criteria such as energy expenditure, fuel consumption, or multi-objective trade-offs between efficiency and safety. Another important direction of research can be related to more realistic agent models. Incorporating inertia, turning-rate limitations and stochastic forcing representing environmental noise would further bridge the gap between theory and practical deployment [colabrese2017flow, alageshan2020machine, piro2022optimal].
@@
From a methodological perspective, the coexistence of ground truth optimal-control solutions and RL-based strategies within the same testbed creates a rare opportunity. It allows quantitative comparison between exact minimum-time trajectories and learned policies in data-driven control. Because the navigation problem is physically transparent, the framework is also well suited for use in teaching, providing an accessible setting in which students can explore nonlinear dynamics, turbulent transport, and modern machine-learning techniques within a unified context.
@@
Overall, our findings demonstrate that RL provides a powerful framework for navigation in complex flow environments, complementing classical optimal control theory and overcoming its primary limitations in chaotic regimes. These results open the door to developing practical autonomous navigation strategies for micro-robots, underwater vehicles, and other active agents operating in realistic oceanic and atmospheric flows.
@@ Acknowledgements.
The work was supported by the Ministry of Science and Higher Education of the Russian Federation and by the Basic research program of HSE University. I am also grateful to the Landau Institute for providing computing resources.
@@ Appendix A Incremental update
Assume for a moment that we are interested in estimating the mean value of a random number generator. At each step, we obtain a single sample, and just before the $n$-th step, our estimate of the mean is given by
$Q_{n}=\frac{r_{1}+r_{2}+\dots+r_{n-1}}{n-1}.$ (55)
@@
where the target represents a potentially noisy indicator of the true value toward which the estimate should move, and $\alpha$ is the step-size or learning rate that controls how strongly the estimate is updated. In the present case, the target corresponds to the new sample $r_{n}$, and the learning rate is $\alpha=1/n$. Because this learning rate decays over time, each update becomes smaller, which guarantees convergence of $Q_{n}$ to the true mean as $n\to\infty$ in accordance with the law of large numbers, see Eq. (55).
@@ Appendix B Policy Gradient Theorem
$\nabla v_{\pi}(s)=\sum_{a}\left[\nabla\pi(a|s)q_{\pi}(s,a)+\pi(a|s)\nabla q_{\pi}(s,a)\right].$ (59)
@@
$\nabla q_{\pi}(s,a)=\sum_{s^{\prime}}p(s^{\prime}|s,a)\nabla v_{\pi}(s^{\prime}),$ (61)
where $p(s^{\prime}|s,a)=\sum_{r}p(s^{\prime},r|s,a)$ denotes the marginal transition probability. Substituting Eq. (61) into Eq. (59) leads to a recursive form,
$\nabla v_{\pi}(s)=\sum_{a}\Big[$ $\nabla\pi(a|s)q_{\pi}(s,a)+$ (62)
$\pi(a|s)\sum_{s^{\prime}}p(s^{\prime}|s,a)\nabla v_{\pi}(s^{\prime})\Big].$
The term involving $\nabla v_{\pi}(s^{\prime})$ can itself be expanded using Eq. (62), and this substitution can be applied recursively to capture contributions from all future states along a trajectory generated by the policy. After the first substitution, we obtain